OpenStudy (anonymous):

Can someone look this over for me?? Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 5, -3, and -1 + 3i f(x) = x4 + 12.5x2 - 50x - 150 f(x) = x4 - 4x3 + 15x2 + 25x + 150 f(x) = x4 - 4x3 - 15x2 - 25x - 150 ---my answer f(x) = x4 - 9x2 - 50x - 150

OpenStudy (anonymous):

@barbillus @SolomonZelman @KlOwNlOvE

OpenStudy (solomonzelman):

A polynomial function with roots a, b, c is written as f(x)=(x-a)(x-b)(x-c) for example, roots -3, 2, 5 give the following f(x) = (x+3)(x-2)(x-5) and to write the full function you just need to expand. Can you write your function and then expand it ?

OpenStudy (solomonzelman):

in your case it is going to be f(x)=(x-a)(x-b)(x-c)(x-d) with roots a b c d

OpenStudy (anonymous):

x4 - 4x3 + 15x2 + 25x + 150 @SolomonZelman

OpenStudy (solomonzelman):

don't think so

OpenStudy (anonymous):

was my first answer right? @SolomonZelman

OpenStudy (anonymous):

?? @SolomonZelman